# Semi-SD: Semi-Supervised Metric Depth Estimation via Surrounding Cameras for Autonomous Driving

### v1

@@ Abstract
In this paper, we introduce Semi-SD, a novel metric depth estimation framework tailored for surrounding cameras equipment in autonomous driving. In this work, the input data consists of adjacent surrounding frames and camera parameters. We propose a unified spatial-temporal-semantic fusion module to construct the visual fused features. Cross-attention components for surrounding cameras and adjacent frames are utilized to focus on metric scale information refinement and temporal feature matching. Building on this, we propose a pose estimation framework using surrounding cameras, their corresponding estimated depths, and extrinsic parameters, which effectively address the scale ambiguity in multi-camera setups. Moreover, semantic world model and monocular depth estimation world model are integrated to supervised the depth estimation, which improve the quality of depth estimation. We evaluate our algorithm on DDAD and nuScenes datasets, and the results demonstrate that our method achieves state-of-the-art performance in terms of surrounding camera based depth estimation quality. The source code will be available on https://github.com/xieyuser/Semi-SD.
Index Terms:
Visual learning, deep learning for visual perception, metric depth estimation, autonomous driving.
@@ I Introduction
Metric depth estimation provides distance perception of the surrounding environment in autonomous driving scenarios, supporting the following tasks such as motion forecasting [2] and path planning [3, 4]. Some existing framworks [5, 6] directly detect depth by introducing LiDAR and Radar sensors, but these approaches lead to increased costs and algorithmic complexity. Currently, autonomous driving perception systems tend to favor visual-only surrounding camera solutions [3, 7] that are easier to implement. Some visual-based depth prediction algorithms [8, 9, 10, 11, 12] use large datasets to train monocular depth estimation (MDE), achieving impressive results in depth estimation as world models. However, the scale ambiguity encountered in monocular depth estimation makes it unsuitable for applications requiring precise depth information in autonomous driving. Surrounding cameras, which inherently provide scale information from extrinsic parameters, have seen rapid development in metric depth prediction [13, 14, 8]. These methods typically use data from the same frame [8] or from the same camera [13, 14], but do not take advantage of the stereo geometric constraints provided by spatial and temporal information, which results in insufficient accuracy and poor generalization. Additionally, some methods [15, 16] use semi-supervised training guidance but face challenges in decoupling complex tasks efficiently. Therefore, the results are not accurate and easy to convergence to local optima. Lastly, these methods fail to integrate semantic information effectively, leading to unclear boundaries in depth maps, particularly in areas where humans rely on semantic cues.
@@
To address these issues, we propose a surrounding camera metric depth estimation framework, named Semi-SD. The results of the paper are briefly presented in Fig. 1. By utilizing two adjacent frames of surrounding camera images, we employ a unified spatial-temporal-semantic transformer [17] to fuse visual features extracted by ResNet [18] and semantic features from a semantic segmentation world model [19]. This fused features are then used for both depth prediction and pose estimation to boost computational efficiency. Furthermore, we integrate depth prediction and the extrinsic parameters of the surrounding camera into the pose estimation network, redesigning a surrounding camera pose estimation module with precise scale information. Additionally, we introduce a curvature loss based on the depth estimation world model [9, 10]. Experimental results demonstrate that the inclusion of this loss function significantly enhances the model’s convergence speed and depth prediction accuracy. Our contributions are summarized as follows:
@@
  • We conduct comprehensive validation on two widely used datasets. The results demonstrate the superior capability of our algorithm to achieve state-of-the-art performance in metric depth estimation for autonomous driving.
@@ III-A Problem Formulation and Semantic Feature Adapter
[…]
@@
(${}_{s}{\mathcal{F}_{(l)}}=DeConvs^{s}_{(l)}\left(Convs^{s}_{(l)}(\mathcal{F}_{(l)})+\mathcal{F}_{seg}\right)$ (1)
@@ III-B Spatial-Temporal-Semantic Transformer
The overall structure of the $l$th STST is shown in Fig. 4. For the feature at the $l$th layer, (${}_{s}\mathcal{F}_{(l)}$, we output the feature (${}_{sts}\mathcal{F}_{(l)}$ with spatial-temporal-semantic information, which retains the same dimension. To reduce computational cost, we first down-sample the feature (${}_{s}\mathcal{F}_{(l)}^{G\times N\times C_{l}\times H_{l}\times W_{l}}$ with different down-sampling ratios for different layers, such that the shape of down-sampled feature matches the unified input shape of STST $G\times N\times\overline{C}\times\overline{H}\times\overline{W}$. Note that $\overline{C}$, $\overline{H}$ and $\overline{W}$ denote the fixed number of channels, feature width, and feature height accepted by STST. Further details are provided in our code.
Then, by applying self-designed surrounding camera attention and adjacent frame attention detailed in Sec. III-B, followed by upsampling to restore the shape, we obtain the desired output (${}_{sts}{\mathcal{F}}_{(l)}^{G\times N\times C_{l}\times H_{l}\times W_{l}}$. Additionally, we use skip connections between the encoder and decoder of the $l$th layer to retain gradient information.
Spatial Surrounding Camera Attention and Temporal Adjacent Frame Attention. By leveraging features from overlapping regions and extrinsic parameters between cameras, we design a spatial attention mechanism for neighboring cameras to inference the metric scale information. Unlike SurroundDepth [16], we reduce computational load by only computing interactions with clockwise neighbors. Additionally, we introduce a temporal attention mechanism for inter-frame feature matching to improve joint pose estimation accuracy. The cross-attention module is shown in Fig. 4, the input feature is (${}_{s}\mathcal{F}_{(l)}^{G\times N\times\overline{C}\times\overline{H}\times\overline{W}}$ and the output feature with same dimension is (${}_{sts}\mathcal{F}_{(l)}^{G\times N\times\overline{C}\times\overline{H}\times\overline{W}}.$
@@
(${}_{s}\mathcal{F}_{(l)}^{G\times N\times\overline{C}\times\overline{H}\times\overline{W}}={{}_{s}\mathcal{F}}_{(l)}^{G\times N\times\overline{C}\times\overline{H}\times\overline{W}}\oplus bc(L_{GN})$ (2)
where $bc$ refers to using the broadcasting mechanism to expand the dimensions of $L_{GN}$, transforming it into $L^{G\times N\times\overline{C}\times\overline{H}\times\overline{W}}_{GN}$. Based on this, we use a linear layer $\left\langle\mathbf{W_{in}},\,\mathbf{b_{in}}\right\rangle$ to map this feature to (${}^{proj}\mathcal{F}_{(l)}=\mathbf{W_{in}}\cdot Flatten(\mathcal{F}_{(l)})+\mathbf{b_{in}}$, where $Flatten$ refers to the operation of converting a tensor into a one-dimensional array. Then, for the $n$th camera, we compute the cross-attention score $A_{N}$ with its adjacent camera. The query, the key and value are defined as
$Q_{n}$ $=_{\quad\,s}^{proj}\mathcal{F}_{(l)}[\,:\,,\,n,\,...]$ (3)
$K_{n+1},\,V_{n+1}$ $=_{\quad\,s}^{proj}\mathcal{F}_{(l)}[\,:\,,\,(n+1)\%N,\,...]$
@@
For the $src$ frame, we compute the cross-attention score $A_{G}$ with its adjacent $tgt$ frame. The query, key and value are defined as defined as
$Q_{src}$ $=_{\quad\,s}^{proj}\mathcal{F}_{(l)}[src,\,...]$ (5)
$K_{tgt},\,V_{tgt}$ $=_{\quad\,s}^{proj}\mathcal{F}_{(l)}[tgt,\,...]$
@@
We then combine the computed features (${}_{attn}\mathcal{F}_{(l)}=\langle A_{N},A_{G}\rangle$ and apply layer normalization followed by linear transformation $\left\langle\mathbf{W_{0}},\,\mathbf{b_{0}}\right\rangle,\left\langle\mathbf{W_{out}},\,\mathbf{b_{out}}\right\rangle$ to project it back to the original dimension:
(${}_{norm}\mathcal{F}_{(l)}$ $=LayerNorm(_{attn}\mathcal{F}_{(l)})$ (7)
(${}_{sts}\mathcal{F}_{(l)}$ $=\mathbf{W_{out}}\cdot ReLU($
@@
III-C Multi-Camera Enhanced Pose Estimation
Unlike most image-based methods [15, 16, 11], we believe that the results of depth map prediction can provide geometric guidance for pose estimation. Furthermore, the coarse-to-fine pose refinement approach used in traditional methods can still be applied in deep neural network. Therefore, we combine the features (${}_{sts}\mathcal{F}_{(l)}$ extracted from the $l$th layer of the STST backbone with the depth prediction (${}_{d}\mathcal{F}_{(l)}$ from the $l$th depth decoder, and then utilize a multi-stage convolutional network to integrate these features.
@@ III-C Multi-Camera Enhanced Pose Estimation
(${}_{p}\mathcal{F}_{(l)}=Convs^{d}_{(l)}(_{sts}\mathcal{F}_{(l)}\otimes\,_{d}\mathcal{F}_{(l)})$ (8)
where $Convs^{d}_{(l)}$ denotes the convolution and sampling kernel. The concatenated features of all $L$ layers will be ${}_{p}\mathcal{F}=\cup_{l=1}^{L}{(_{p}\mathcal{F}_{(l)})}.$ After aligning the feature dimensions, a lightweight pose decoder with convolution and mean operations is employed to generate the final joint pose estimation $\mathcal{P}_{s\to t}$. The overview of pose decoder and $l$th depth decoder are shown in Figure 5.
@@ III-D Implementation Details
In our framework, we predict a full-size depth map, and then reshape it to the $k$th level ($30<k\leq K,K=$ 3 in our experiments) in a pyramid structure. The loss is calculated across all $K$ levels.
@@
We define a $t$-step ($30<t\leq T,T=$ 3 in our experiments) depth gradient operator for coordinate $x,y$ as
@@
Reprojection Loss and Semantic Loss. We predict the inverse depth $\mathcal{D}^{src}_{(n)(k)}$ of $n$th image in $k$th level and relative pose $\mathcal{P}_{n}$. The operator $\Lambda_{(n)}^{(k)}$ is defined as reprojecting image ($\mathcal{I}^{src}_{(n)(k)}$ to 00th layer ($\mathcal{I}^{tgt}_{(n)(0)}$ by combining with sky and optional vehicle body occlusion mask (($\mathcal{M}^{src}_{(n)(k)}$, ($\mathcal{M}^{tgt}_{(n)(k)}$):
@@
where $\pi^{src}_{(n)(k)}$ and $\pi^{tgt}_{(n)(0)}$ are the intrinsic of ($\mathcal{I}^{src}_{(n)(k)}$ and ($\mathcal{I}^{tgt}_{(n)(0)}$, respectively. $\odot$ denotes the Hadamard product applied element-wisely to the predicted depth. We calculate reprojection loss $\mathcal{L}_{rep}$ via L1 distance and SSIM distance [33] between $N$ $src$ frames ($\mathcal{I}^{src}_{(n)(k)}$ and $N$ $tgt$ frames (${}^{M}\mathcal{I}^{tgt}_{(n)(0)}\leftarrow M^{tgt}_{(n)(0)}\odot\mathcal{I}^{tgt}_{(n)(0)}$. $\mathcal{L}_{rep}$ is defined as
@@
As for semantic loss, a pre-trained semantic world model $Seg$ [32] is used to extract semantic features ($\mathcal{I}^{tgt}_{(0)}$. The goal is to align these features in a high-dimensional space during the training process. The alignment is supervised by minimizing
@@ IV Experiments
Ours-Medium 800 $\times$ 448 Semi 0.167 2.686 11.407 0.283 0.792 0.904 0.953
@@
Ours-Medium 800 $\times$ 448 Semi 0.197 2.624 6.094 0.297 0.789 0.903 0.946
@@ IV-A Environmental Setup
Baselines and Metrics. We compare our results with R3D3 [13], VFDepth [29], scale-aware SurroundDepth [16], and M2-Depth [15]. It is pertinent to note that both SurroundDepth [16] and M2-Depth [15] utilize the structure-from-motion (SfM) method to generate sparse depth for training supervision, contradicting their claim of being self-supervised in their papers. Therefore, in the comparative experiments of this paper, we categorize them as semi-supervised methods. Following prior works [35, 15, 16], the evaluation metrics we used are $Abs.Rel.$, $Sq.Rel.$, $RMSE$, $RMSE\,log$, and $\delta$.
Datasets. We train and evaluate our framework on two public datasets, including DDAD [1] and nuScenes [34]. For the DDAD [1] dataset, we first crop the image from the top-left corner to a size of 1936$\times$1084, and then resize it to 800$\times$448. For the nuScenes [34] dataset, we directly resize the images to 800$\times$448. To the best of our knowledge, the image sizes employed in our work are the largest, which helps improve the extraction of detailed features and their subsequent refinement.

### v3

# Semi-SMD: Semi-Supervised Metric Depth Estimation via Surrounding Cameras for Autonomous Driving

@@ Abstract
In this paper, we introduce Semi-SMD, a novel metric depth estimation framework tailored for surrounding cameras equipment in autonomous driving. In this work, the input data consists of adjacent surrounding frames and camera parameters. We propose a unified spatial-temporal-semantic fusion module to construct the visual fused features. Cross-attention components for surrounding cameras and adjacent frames are utilized to focus on metric scale information refinement and temporal feature matching. Building on this, we propose a pose estimation framework using surrounding cameras, their corresponding estimated depths, and extrinsic parameters, which effectively address the scale ambiguity in multi-camera setups. Moreover, semantic world model and monocular depth estimation world model are integrated to supervise the depth estimation, which improve the quality of depth estimation. We evaluate our algorithm on DDAD and nuScenes datasets, and the results demonstrate that our method achieves state-of-the-art performance in terms of surrounding camera based depth estimation quality. The source code is available on GitHub [^1] https://github.com/xieyuser/Semi-SMD.
@@ I Introduction
Metric depth estimation provides absolute distance perception of the surrounding environment in autonomous driving scenarios, supporting the following tasks such as motion forecasting [1] and path planning [2, 3]. Some existing framworks [4, 5] directly detect depth by introducing LiDAR and Radar sensors, but these approaches lead to increased costs and algorithmic complexity. Currently, autonomous driving perception systems tend to favor visual-only surrounding camera solutions [2, 6] that are easier to implement. Some visual-based depth prediction algorithms [7, 8, 9, 10, 11] use large datasets to train monocular depth estimation (MDE), achieving impressive results in depth estimation as world models. However, the scale ambiguity encountered in monocular depth estimation makes it unsuitable for applications requiring precise depth information in autonomous driving. Surrounding cameras, which inherently provide scale information from extrinsic parameters, have seen rapid development in metric depth prediction [12, 13, 7]. These methods typically use data from the same frame [7] or from the same camera [12, 13], but do not take advantage of the stereo geometric constraints provided by spatial and temporal information, which results in insufficient accuracy and poor generalization. Additionally, some methods [14, 15] use semi-supervised training guidance but face challenges in decoupling complex tasks efficiently. Therefore, the results are not accurate and easy to convergence to local optima. Lastly, these methods fail to integrate semantic information effectively, leading to unclear boundaries in depth maps, particularly in areas where humans rely on semantic cues.
@@
To address these issues, we propose a surrounding camera metric depth estimation framework, named Semi-SMD. The results of the paper are briefly presented in Fig. 1. By utilizing two adjacent frames of surrounding camera images, we employ a unified spatial-temporal-semantic transformer [17] to fuse visual features extracted by ResNet [18] and semantic features from a semantic segmentation world model [19]. This fused features are then used for both depth prediction and pose estimation to boost computational efficiency. Furthermore, we integrate depth prediction and the extrinsic parameters of the surrounding camera into the pose estimation network, redesigning a surrounding camera pose estimation module with precise scale information. Additionally, we introduce a curvature loss based on the depth estimation world model [8, 9]. Experimental results demonstrate that the inclusion of this loss function significantly enhances the model’s convergence speed and depth prediction accuracy. Our contributions are summarized as follows:
@@
  • We conduct comprehensive validation on two widely used datasets. The results demonstrate the superior capability of our algorithm to achieve SOTA performance in metric depth estimation for autonomous driving.
@@ III-A Problem Formulation and Semantic Feature Adapter
A brief overview of our framework is illustrated in Fig. 2. We extract visual features from combined images $\mathcal{I}^{G\times N\times 3\times H\times W}$ by ResNet [18], and the output of the $L$-layers extracted features is denoted as $\mathcal{F}_{vis}=\{\mathcal{F}_{(l)}^{G\times N\times C_{l}\times H_{l}\times W_{l}},\,0\leq l<L\}$, where $C_{l}$, $H_{l}$, and $W_{l}$ represent the number of channels, feature height, and feature width at the $l$th layer, respectively. Meanwhile, semantic features $\mathcal{F}_{seg}$ are obtained from combined images $\mathcal{I}^{G\times N\times 3\times H\times W}$ by the frozen SAM feature encoder [31]. Features $\mathcal{F}_{vis}$ and $\mathcal{F}_{seg}$ are then passed through a self-designed spatial-temporal-semantic transformer (STST) module (detailed in Sec. III-B) to fuse features. This fused feature, which aggregates temporal, spatial, and semantic information, serves as the input for both the depth decoder and the pose decoder, thereby avoiding redundant feature extraction [15, 14] from the raw images and making the network more lightweight. In Sec. III-C, we introduce the structure of the STST-enhanced joint pose estimation network. Following this, we detail the implementation of our loss function and network in Sec. III-D.
@@
$_{s}{\mathcal{F}_{(l)}}=DeConvs^{s}_{(l)}\left(Convs^{s}_{(l)}(\mathcal{F}_{(l)})+\mathcal{F}_{seg}\right)$ (1)
@@ III-B Spatial-Temporal-Semantic Transformer
The overall structure of the $l$th STST is shown in Fig. 4. For the feature at the $l$th layer, $_{s}\mathcal{F}_{(l)}$, we output the feature $_{sts}\mathcal{F}_{(l)}$ with spatial-temporal-semantic information, which retains the same dimension. To reduce computational cost, we first down-sample the feature $_{s}\mathcal{F}_{(l)}^{G\times N\times C_{l}\times H_{l}\times W_{l}}$ with different down-sampling ratios for different layers, such that the shape of down-sampled feature matches the unified input shape of STST $G\times N\times\overline{C}\times\overline{H}\times\overline{W}$. Note that $\overline{C}$, $\overline{H}$ and $\overline{W}$ denote the fixed number of channels, feature width, and feature height accepted by STST. Further details are provided in our code.
Then, by applying self-designed surrounding camera attention and adjacent frame attention detailed in Sec. III-B, followed by upsampling to restore the shape, we obtain the desired output $_{sts}{\mathcal{F}}_{(l)}^{G\times N\times C_{l}\times H_{l}\times W_{l}}$. Additionally, we use skip connections between the encoder and decoder of the $l$th layer to retain gradient information.
Spatial Surrounding Camera Attention and Temporal Adjacent Frame Attention. By leveraging features from overlapping regions and extrinsic parameters between cameras, we design a spatial attention mechanism for neighboring cameras to inference the metric scale information. Unlike SurroundDepth [15], we reduce computational load by only computing interactions with clockwise neighbors. Additionally, we introduce a temporal attention mechanism for inter-frame feature matching to improve joint pose estimation accuracy. The cross-attention module is shown in Fig. 4, the input feature is $_{s}\mathcal{F}_{(l)}^{G\times N\times\overline{C}\times\overline{H}\times\overline{W}}$ and the output feature with same dimension is $_{sts}\mathcal{F}_{(l)}^{G\times N\times\overline{C}\times\overline{H}\times\overline{W}}.$
@@
$_{s}\mathcal{F}_{(l)}^{G\times N\times\overline{C}\times\overline{H}\times\overline{W}}={{}_{s}\mathcal{F}}_{(l)}^{G\times N\times\overline{C}\times\overline{H}\times\overline{W}}\oplus bc(L_{GN})$ (2)
where $bc$ refers to using the broadcasting mechanism to expand the dimensions of $L_{GN}$, transforming it into $L^{G\times N\times\overline{C}\times\overline{H}\times\overline{W}}_{GN}$. Based on this, we use a linear layer $\left\langle\mathbf{W_{in}},\,\mathbf{b_{in}}\right\rangle$ to map this feature to $^{proj}\mathcal{F}_{(l)}=\mathbf{W_{in}}\cdot Flatten(\mathcal{F}_{(l)})+\mathbf{b_{in}}$, where $Flatten$ refers to the operation of converting a tensor into a one-dimensional array. Then, for the $n$th camera, we compute the cross-attention score $A_{N}$ with its adjacent camera. The query, the key and value are defined as
$Q_{n}$ $=_{\quad\,s}^{proj}\mathcal{F}_{(l)}[\,:\,,\,n,\,$.] (3)
$K_{n+1},\,V_{n+1}$ $=_{\quad\,s}^{proj}\mathcal{F}_{(l)}[\,:\,,\,(n+1)\%N,\,$.]
@@
For the srcsrc frame, we compute the cross-attention score $A_{G}$ with its adjacent tgttgt frame. The query, key and value are defined as defined as
$Q_{src}$ $=_{\quad\,s}^{proj}\mathcal{F}_{(l)}[src,\,$.] (5)
$K_{tgt},\,V_{tgt}$ $=_{\quad\,s}^{proj}\mathcal{F}_{(l)}[tgt,\,$.]
@@
We then combine the computed features $_{attn}\mathcal{F}_{(l)}=\langle A_{N},A_{G}\rangle$ and apply layer normalization followed by linear transformation $\left\langle\mathbf{W_{0}},\,\mathbf{b_{0}}\right\rangle,\left\langle\mathbf{W_{out}},\,\mathbf{b_{out}}\right\rangle$ to project it back to the original dimension:
${}_{norm}\mathcal{F}_{(l)}$ $=LayerNorm(_{attn}\mathcal{F}_{(l)})$ (7)
${}_{sts}\mathcal{F}_{(l)}$ $=\mathbf{W_{out}}\cdot ReLU($
@@
III-C Multi-Camera Enhanced Pose Estimation
Unlike most image-based methods [14, 15, 10], we believe that the results of depth map prediction can provide geometric guidance for pose estimation. Furthermore, the coarse-to-fine pose refinement approach used in traditional methods can still be applied in deep neural network. Therefore, we combine the features $_{sts}\mathcal{F}_{(l)}$ extracted from the $l$th layer of the STST backbone with the depth prediction $_{d}\mathcal{F}_{(l)}$ from the $l$th depth decoder, and then utilize a multi-stage convolutional network to integrate these features.
@@ III-C Multi-Camera Enhanced Pose Estimation
$_{p}\mathcal{F}_{(l)}=Convs^{d}_{(l)}(_{sts}\mathcal{F}_{(l)}\otimes\,_{d}\mathcal{F}_{(l)})$ (8)
where $Convs^{d}_{(l)}$ denotes the convolution and sampling kernel. The concatenated features of all $L$ layers will be $_{p}\mathcal{F}=\cup_{l=1}^{L}{(_{p}\mathcal{F}_{(l)})}.$ After aligning the feature dimensions, a lightweight pose decoder with convolution and mean operations is employed to generate the final joint pose estimation $\mathcal{P}_{s\to t}$. The overview of pose decoder and $l$th depth decoder are shown in  Fig. 5.
@@ III-D Implementation Details
In our framework, we predict a full-size depth map, and then reshape it to the $k$th level ($0<k\leq K,K=3$ in our experiments) in a pyramid structure. The loss is calculated across all $K$ levels.
@@
We define a $t$-step ($0<t\leq T,T=3$ in our experiments) depth gradient operator for coordinate $x,y$ as
@@
Reprojection Loss and Semantic Loss. We predict the inverse depth $\mathcal{D}^{src}_{(n)(k)}$ of $n$th image in $k$th level and relative pose $\mathcal{P}_{n}$. The operator $\Lambda_{(n)}^{(k)}$ is defined as reprojecting image $\mathcal{I}^{src}_{(n)(k)}$ to 0th layer $\mathcal{I}^{tgt}_{(n)(0)}$ by combining with sky and optional vehicle body occlusion mask ($\mathcal{M}^{src}_{(n)(k)}$, $\mathcal{M}^{tgt}_{(n)(k)}$):
@@
where $\pi^{src}_{(n)(k)}$ and $\pi^{tgt}_{(n)(0)}$ are the intrinsic of $\mathcal{I}^{src}_{(n)(k)}$ and $\mathcal{I}^{tgt}_{(n)(0)}$, respectively. $\odot$ denotes the Hadamard product applied element-wisely to the predicted depth. We calculate reprojection loss $\mathcal{L}_{rep}$ via L1 distance and SSIM distance [32] between NN srcsrc frames $\mathcal{I}^{src}_{(n)(k)}$ and NN tgttgt frames $^{M}\mathcal{I}^{tgt}_{(n)(0)}\leftarrow M^{tgt}_{(n)(0)}\odot\mathcal{I}^{tgt}_{(n)(0)}$. $\mathcal{L}_{rep}$ is defined as
@@
As for semantic loss, a pre-trained semantic world model $Seg$ [31] is used to extract semantic features $\mathcal{I}^{tgt}_{(0)}$. The goal is to align these features in a high-dimensional space during the training process. The alignment is supervised by minimizing
@@ IV Experiments
Ours 800 $\times$ 448 Semi 0.167 2.686 11.407 0.283 0.792 0.904 0.953
@@
Ours 800 $\times$ 448 Semi 0.197 2.624 6.094 0.297 0.789 0.903 0.946
@@ IV-A Environmental Setup
Baselines and Metrics. We compare our results with R3D3 [12], VFDepth [28], scale-aware SurroundDepth [15], and M2-Depth [14]. It is pertinent to note that both SurroundDepth [15] and M2-Depth [14] utilize the structure-from-motion (SfM) method to generate sparse depth for training supervision, contradicting their claim of being self-supervised in their papers. Therefore, in the comparative experiments of this paper, we categorize them as semi-supervised methods. Following prior works [34, 14, 15], the evaluation metrics we used are $Abs.Rel.$, $Sq.Rel.$, RMSERMSE, RMSElogRMSE\,log, and $\delta$.
[…]